Steve has 36 coins consisting of dimes and nickels. If the value of the coins is $3.10, then how many dimes does he have?
Answer by mukhopadhyay(490)

dimes = x
nickels = y
x+y = 36
=> x = 36-y;
10x + 5y = 310
=> 10(36-y) + 5y = 310
=> 360-10y+5y = 310
=> 5y = 50
=> y = 10;
x = 36-y
=> x = 26
Answer: Steve has 26 dimes and 10 nickels